OpenStudy (jewlzme17):

Which Symmetry? Options: Y axis symmetry,X axis symmetry, and Origin symmetry. 1. x-y^2=0 2. y= x^4-x^2+3 3.y=x/x^2+1 4. y=(sqrt 9-x^2) 5.xy^2+10=0 6. xy=4

OpenStudy (jewlzme17):

Please Explain how you did it!

OpenStudy (anonymous):

See if this helps. ODD Functions: Opposite X's and Opposite Y's and symmetrical to the origin. EVEN Functions: Opposite X's nut Equal Y's. In that case even functioms would be symmetrical to the y axis. If you have opposite x's and opposite y's then that means you are symmetrical to the origin. Ex: (5,10) (-5,-10). With X- Axis, your Y values change. Ex: (-3,5) (-3,-5)

OpenStudy (anonymous):

Try and solve the first one, and ill let you know if you got it.

OpenStudy (jewlzme17):

okay so x-y^2=0 will go to (-x)-y^2=0 so it would be y axis symm

OpenStudy (anonymous):

Let me ask this first. Do you agree at x^2, it creates a parabla

OpenStudy (jewlzme17):

yes

OpenStudy (anonymous):

The equation ultimately ends up being. X=Y^2. You probabably recognize the opposite of that, Y=X^2. With Y=x^2 the parabla would make a U shape that would mean the ordered pairs would have opposite x's but equal y's or in other words, symmetrical to the y axis.. Now if X is what your trying to find, what must it be symmetrical to?

OpenStudy (anonymous):

Does any of that make sense to you? Do i need to clarify anything?

OpenStudy (jewlzme17):

yea it kind of does! ha but for the x axis you change the y. So from (x,y) to (x,-y)

OpenStudy (anonymous):

Yes exactly. Think of the parabola on it's side. Like how Y=x^2 the parabola went through (0,0) and the two pieces curved up on each side of the y axis. In this case, your just flipped on the x-axis, and instead of your x values changing, your y values change, exactly like you mentioned

OpenStudy (jewlzme17):

okay so it would be x-y^2=0 to x+y^2=0

OpenStudy (anonymous):

No. The question you asked, was what would be the symmetry. The symmetry would be over the x-axis. x-y^2=0 to x+y^2=0 can lead to 2 separate functions: x-y^2=0 ends up being x=y^2. x+y^2=0 ends up being x=-y^2. Making the parabola negative negates it, and flips facing the other direction.

OpenStudy (jewlzme17):

would 5 be xy^2+10=0 to x(-y)^2+10=0 to xy^2+10=0 so it is the x -axis? But no I know that it's just I'm showing my work on how.

OpenStudy (anonymous):

Is your question if that your work is right?

OpenStudy (jewlzme17):

yea! and if the axis is correct

OpenStudy (anonymous):

Your axis is correct. It is symmetrical over the x-axis. However im confused with what you typed "xy^2+10=0 to x(-y)^2+10=0 to xy^2+10=0".

OpenStudy (anonymous):

What I would do personally, is make a table with your x values. and just plug them in and see if they follow any particular rule.

OpenStudy (anonymous):

Heres an example.

OpenStudy (jewlzme17):

that was me changing symbols and then solving ha idk that is what my teacher taught me.

OpenStudy (anonymous):